# Operator World Models for Reinforcement Learning

### v1

@@ C.3 Convergence Rates for POWR
Since the setting of Cor. 6 implies that ${P}_{t}$ are $({\mathcal{G}},{\mathcal{F}})$-compatible for all $t$, and Asm. 2 is holding, then $q({P}_{t})$ and $q_{n}({P}_{t})$ belong to ${\mathcal{G}}$ for all $({P}_{t})_{t\in\mathbb{N}}$. This assures that we can use the statistical learning bounds Proposition C.2 into Thm. 7, yielding the final bound. ∎
@@ D.1 Additional Results
In this section, we delve deeper into the empirical outcomes of our methodology. We present a boxplot that shows the average timestep at which a reward threshold is met during the training phase. The testing environments are the same as introduced previously, with reward thresholds being the standard ones given in [36], except for the Taxi-v3 environment, where it is marginally lower. Interestingly, in this environment, only DQN and our algorithm are capable of achieving the original threshold within $1.5\times 10^{6}$ timesteps during the training. On the other hand, the new lower threshold is also reached by the PPO algorithm.
As depicted in Fig. 2, our approach can attain the desired reward quicker than the competing algorithms. Furthermore, when considering our method, the timestep at which the threshold is reached exhibits a lower variance than other techniques. This implies that our approach requires a stable amount of timesteps to learn how to solve a specific environment.
@@
Figure 2: This boxplot visually represents the mean timestep at which various algorithms attain a specified reward threshold during their training. The reward targets are set at $0.8$ for FrozenLake-v1, $6$ for Taxi-v3, and $-110$ for MountainCar-v0. The absence of a box indicates that the corresponding algorithm was unable to meet the reward threshold within the training process.
D.2 Hyperparameters
To test our approach, we fine-tuned the algorithm through a grid search. Table 1 reports the optimal parameters for each testing environment. Here, t_epochs denotes the training epochs used for data collection t_epochs before updating the policy. Then, n_iter_pmd refers to the number of policy mirror descent iterations to update our policy.
Parameter FrozenLake-v1 Taxi-v3 MountainCar-v0
$\eta$ 1 1 0.1
$\lambda$ $10^{-6}$ $10^{-6}$ $10^{-6}$
$\gamma$ 0.99 0.99 0.99
t_epochs 5 5 1
n_iter_pmd 10 10 10
Table 1: Hyperparameters for using our method in the proposed environments.
D.3 Other methods
We compare the performance of our algorithm with several baselines. In particular, we considered A2C [37], DQN [4], TRPO [7] and PPO [6], which we implemented using the stable baselines library [38]. In the end, we used the standard hyperparameters in [38].

### v2

@@ C.3 Convergence Rates for POWR
Since the setting of Cor. 6 implies that $\mathsf{P}_{t}$ are $({\mathcal{G}},{\mathcal{F}})$-compatible for all $t$, and Asm. 2 is holding, then $q(\mathsf{P}_{t})$ and $q_{n}(\mathsf{P}_{t})$ belong to ${\mathcal{G}}$ for all $(\mathsf{P}_{t})_{t\in{\mathbb{N}}}$. This assures that we can use the statistical learning bounds Proposition C.2 into Thm. 7, yielding the final bound. ∎
@@ D.1 Additional Results
In Fig. 2 we show the average timestep at which a reward threshold is met during the training phase. The testing environments are the same as introduced previously, with reward thresholds being the standard ones given in [25], except for the Taxi-v3 environment, where it is marginally lower. Interestingly, in this environment, only DQN and our algorithm are capable of achieving the original threshold within $1.5\times 10^{6}$ timesteps during the training. On the other hand, the new lower threshold is also reached by the PPO algorithm.
Our approach attain the desired reward quicker than the competing algorithms. Furthermore, the timestep at which POWRreaches the threshold exhibits a lower variance compared to other techniques. This implies that our approach requires a stable amount of timesteps to learn how to solve a specific environment.
@@
Figure 2: Mean timestep at which various algorithms attain a specified reward threshold during their training. The reward targets are set at $0.8$ for FrozenLake-v1, $6$ for Taxi-v3, and $-110$ for MountainCar-v0. The absence of a box indicates that the corresponding algorithm was unable to meet the reward threshold within the training process.
D.2 Other methods
We compare the performance of our algorithm with several baselines. In particular, we considered A2C [40], DQN [4], TRPO [7] and PPO [6], which we implemented using the stable baselines library [47]. We used the standard hyperparameters in [47].